Mrs Daniels has a box of craft sticks for her class.if she gives each student 8 craft sticks she will have 4 craft sticks left.if she gives only 5 craft sticks to each student she will have 40 craft sticks left.how many students are there in the class?

2) Ellen spent 20% of her money on a dress. She spent 2/5 of the remainder on a pair of shoes.she had $72 left.how much money did ellen have at first?

Answer by ikleyn(52792)

.

Solution to Problem 1.


Let n be the number of students in the class.


Then from the condition, you have this equation


    8n + 4 = 5n + 40.


Now solve it


    8n - 5n = 40 - 4

    3n      = 36

     n      = 36/3 = 12.    ANSWER


Answer.  There are 12 students in the class.

Solved.


/\/\/\/\/\/\/\/

Hello, there is a RULE, a POLICY and a REQUIREMENT at this forum:

    ONE and ONLY ONE problem/question per post.


Please, respect the tutors --- do not violate this rule.

Those who violate it, work AGAINST their own interests.


Have a nice day (!)